A Bibliography of Publications in Supercomputer

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

25 October 2010 Version 1.10

Title word cross-reference

1 [Lt93]. 2 [DHRW92]. 3 [RNHW90]. R^{3} [MC96]. y′′ = f(t, y) [CPS97].

-d [RNHW90, DHRW92, Lt93].

1 [BHW^{+}93, Ess88, McB91, McB93]. 1-VX [Ess88]. 10P [Van89]. 1100/90 [vv88].

2 [BW90, LtW88, LJ88, LCH^{+}88, LC97, Pin94, WTW90a, WTW90b]. 2-D [WTW90a, WTW90b]. 200 [MH92]. 205 [LtW88, Van89].

3 [CS90, EM91, HG89, TM90, de 89]. 3-D [CS90, HG89, TM90, de 89]. 3/44 [MSW92b]. 3090 [HG89, RRS88, vv88]. 3090/VF [GKMR89, HG89].

4 [NMW93, de 89]. 432 [MS90a]. 44 [MSW92b].

500 [SNtW94].

6000 [DMR91].

77 [Lan88].

860 [McB91]. 8X [Rei87, LCH^{+}88].

90 [MSW92a, vv88]. ’91 [Ano91a]. 990 [vv88].

Abstract [Gar90]. academic [AKB^{+}97]. access [Ruh90, VD92a]. address [ACK94]. advanced [BP97]. air [BHWZ96, ZWMV92]. algebra [DMR91, DD91, RRS88]. algebraic [RNHW90]. algorithm [LtW88, MC96, OYI91, van93a]. algorithms [CM88, HS90, Hem84, KNG^{+}93, MM89, MS89, Ove90, PL90, Roy92, ZWS89, van91]. Alliant [BPA90, de 89]. alloys [CLHP93]. Alternative [MST96]. Amdahl [DHW92, SS95, WCMJ89]. Amdahl/Fujitsu [DHW92]. AMT [Coo90]. Analysis [DS96, DS97, NAW^{+}96, Sim95, ACM^{+}91, ASN^{+}91, Coi90, Hoc92, PHH91, Roy92, ZPN95, van91]. Application [GK95, KOS^{+}91, Sol89a, NMW93, Pud89, Van92a]. applications [BP97, DVB^{+}97, MB90, O’N88]. approach [Emm89, IG91]. architecture [AHJS90, BDB93, HMN^{+}92]. Architectures [Van95a, Fis90, KC95, Van96, Van97]. arithmetic [BP93]. artificial [Haa89]. aspects [CV93, Van88a, Van88b]. Asynchronous [ALM96, MS89, TP93]. atmospheric [Got92, JH90, Pud89]. Australian [Bar88]. automatic [LCH^{+}88]. automatically [Nob89].

based [BDB93, LJ88, MEGS97, OYI91, ZPN95]. basic [GW91]. Basis [DS96, DS97, Sim95]. behavior [FZM91, Pan90]. Belgian [Van92a]. Benard [TVV96]. Benchmark [Dec88, DH95, FGHv89, Gen89, Haa89, Hoc92, Was93, WGOY91, vd92b, Van93b, Van95b, van93c]. Benchmarking [UT91, KSS96]. benchmarks [HH96, KOS^{+}91, OP93, SH95, SS95, Lan88]. beryllium [REKP94]. Bézier [CM93]. bi [GS90, FZM91]. Bi-CG [FZM91]. bi-directional [GS90]. binary [CLHP93, Dow92, OYI91]. biology [Wit91]. Bitonic [Ove92a]. Black [van93a]. BLAS [WCMJ89]. blending [MS92]. Block [RRS88, PS97, Tsu91, Zha89]. block-structured [Tsu91]. Blocks [MS96]. Boolean [OYI91]. bottleneck [ZPN95]. bound [PL90]. boundary [SSL88]. branch [PL90]. branch-and-bound [PL90]. Building [MS96]. bulk [REKP94]. Bull [vv88]. Butterfly [O’N88].

C [MSW92a]. C-90 [MSW92a]. cache [Get95]. Calculation [Pin94, Mur91]. calculations [MSW92a, MSW92b, WM91]. Calgary [BW90]. Caltech [MB88]. careful [SH95]. cavity [TM90]. ccMBPT [MSW92a, MSW92b, WM91]. cell [DNT96]. center [Ben89]. CERFACS [Ben89]. CG [FZM91]. chain [TP93]. challenge [Oak90]. change [MS95]. Chaosynth [Mir94]. Characterisation [ZPN95, Get95]. characteristics [NMW93]. chemical [NMW93]. chemistry [GKMR89]. Choleski [Mod89]. Cholesky [LC97]. circuit [WE91]. circulation [RD92]. CIT [SRS^{+}90]. Cluster [MST96, REKP94, SH97]. clusters [LCH^{+}96]. CM [MH92]. CM-200 [MH92]. coast [Van92a]. code [ACM^{+}91, CPS97, Paj90]. codes [Dec88, PS92]. Collection [DH95]. combinatorial [PL90]. comfort [TP89]. Communication [CV93, HH96, MH92]. comparative [LCH^{+}88, Nob89]. Comparing [BHWZ96]. Comparison [DHW92, LCH^{+}96, McB91]. compiler [AHJS90, Tsu91]. compilers [Nob89]. complex [MDL^{+}89]. computation [Ben89, Gar90, NK90, SDV89, WM91]. Computational [Got91, Hoc95, GKMR89, de 89]. computer [BW90, BHW^{+}93, BPA90, DNT96, DHW92, Emm89, MDL^{+}89, Mir94, MSW92a, MSW92b, SDV89]. Computers [DV92, BHWZ96, McB91, Van90b]. Computing [DS96, DS97, EH95, Har96, Har97, MST96, Sim95, SDMS97, LCHJ92, LCH^{+}96, MB90, MS95, Mir94, MS90b]. concept [Tro89]. Concurrent [GSv90, WM91, MB88]. condition [Pin94]. confined [ACM^{+}91]. conjugate [CV93, ZWS89]. connectivity [RTY90]. considerations [LHHJ91, Sv96]. construct [MC96]. content [SS95]. convection [CS90, TM90, TVV96]. convergence [FZM91]. Convexity [MM89]. convolution [Ess88]. coprocessor [Lt93]. correlation [MSW92a, MSW92b, WM91]. CPU [Lan88]. CPU-benchmarks [Lan88]. Cray [MS90a, PS97, RNHW90, SNtW94, WTW90a, WTW90b]. crunchers [WGOY91]. Current [Kha91]. curve [MC96]. curves [CM93]. Cyber [LtW88, Van89, vv88].

d [RNHW90, CS90, DHRW92, HG89, Lt93, TM90, WTW90a, WTW90b, de 89]. DAP [Coo90]. Daresbury [GSv90]. data [MC96, SNtW94]. debugger [GH89]. decision [OYI91]. decompilation [OP93]. decomposition [Lt93, MS90a, Pin92]. degree [Haa89]. Denmark [ACH^{+}89]. dependent [TVV96]. Description [Van95a, Van96, Van97]. Design [Tsu91, Coi90]. determinant [Mur91]. Developments [HS96, Sch97]. device [Lt93]. diagrams [OYI91]. different [HS90, Lan88]. digitized [MM89]. dimensional [SRS^{+}90]. direction [Van95b, van93c]. directional [GS90]. discharges [WTW90a, WTW90b]. discovery [Ric91]. discrete [DHRW92]. dispersion [Got92]. distance [MH92]. distributed [ASN^{+}91, CV93, KNG^{+}93, PdR91, PZN96]. distributed-memory [CV93, KNG^{+}93]. Domain [Pin92, Lt93, MS90a]. down [Bar88]. DPS [vv88]. DPS-90 [vv88]. drug [Ric91]. dynamics [BDB93, KNG^{+}93].

economic [NK90]. eddy [BPA90]. effect [Get95]. Effective [Roy92]. effects [Zha89]. efficient [CK92, GW91, Ove91b, Paj90]. effort [FGHv89]. eigenproblems [PS92]. electron [DHRW92, MSW92a, MSW92b, WM91]. element [TM90, Zoi87]. embedded [DVB^{+}97]. engineering [Got91]. enhancement [LHHJ91]. Environment [DV92, ACK94, ASN^{+}91, Buz90, Pin92, SH97]. equation [GS90, SSL88]. Equations [DV92, McB93, Zil96]. essential [SH95]. estimating [DHNP95]. Eta [Van89]. Eta-10P [Van89]. EuroBen [Van93b, Van95b, van92b, vd92b, van93c]. Europe [EH95, Har96, Har97, Oak90]. European [Ben89, FGHv89]. evaluation [BP93, LJ88, NH91, Ove91a]. evolution [MDL^{+}89]. Experience [BHW^{+}93, GH89]. Experiences [Van90b, BW90, VD92a, Vul93]. experimental [GH89]. Experiments [ASN^{+}91]. explicit [LC97]. Exploitation [Rei87].

facilities [Bar88, MB88]. factoring [LtW88]. factorizations [LC97]. factors [Mod89]. family [Ove90]. Fast [Ruh90, WE91, WH90]. Features [Rei87]. featuring [Van96, Van97]. FFT [DHW92, van91]. field [IG91]. finite [TM90, Zoi87]. finite-element [TM90]. Finland [Nie90]. fire [IG91]. First [BW90, HS90, Ove91b, REKP94]. first-order [Ove91b]. five [Pan91]. five-year [Pan91]. Flood [RTY90]. Flosolver [SDV89]. Flow [SDV89]. flows [de 89]. forecast [Pan91]. forecasting [PSNK96]. Fortran [AHJS90, DV92, Lan88, LCH^{+}88, LC97, Nob89, Paj90, RTY90, Rei87, Sol89b, Sol89c]. Fourier [WH90]. framework [Hoc92]. Fujitsu [DNT96, DHW92]. Fukuoka [Ano91b]. fully [Ove90]. fully-vectorizable [Ove90]. functional [MC96]. functions [NH91, OYI91]. Further [Vul93]. fusion [ACM^{+}91]. future [Fis90]. FX [BPA90, de 89]. FX/4 [de 89].

galaxies [HB90]. general [Ove91b, RD92].

Generalized [van93a]. Generation [Paj90, Gen88, IK90, Kle89, OYWK91, van90a]. geometry [SRS^{+}90]. Givens [PZN96]. Global [PSNK96, ACK94, JH90]. GMRESR [Vul93]. GP1000 [O’N88]. gradient [CV93, ZWS89]. Grenoble [ALM96]. gridded [MC96]. grids [Pin92].

heat [CS90]. High [DS96, DS97, EH95, Har96, Har97, MDL^{+}89, Sim95, SDMS97, Ben89, DD91, MS95, Mir94, LC97]. High-Performance [SDMS97, DD91, Mir94]. Highly [GW91]. Hitachi [Van93b]. Honeywell [vv88]. Honeywell-Bull [vv88]. HPC [DVB^{+}97, MS96, MS97]. HPCI [AKB^{+}97]. HPF [BP97]. Hypercube [CM88, Ove92b, MM89, Roy92, SRS^{+}90].

I/O [LHHJ91, VD92a]. i860 [BDB93]. IBM [DMR91, GKMR89, HG89, LC97, NMW93, RRS88, vv88]. image [OYWK91, SRS^{+}90]. implementation [CK92, MS90a, Tsu91, TVV96, Zil93, van93a]. Implementing [WCMJ89]. independent [Emm89]. indirect [UT91]. Individual [OP93]. Industrial [DVB^{+}97, MS97, AKB^{+}97]. Industry [SDMS97]. inevitable [Fis90]. information [SH95]. inhomogeneous [GS90, Hv90]. Institute [ALM96]. instruction [HMN^{+}92]. instructions [UT91]. integral [SSL88]. Integrated [Coi90]. Intel [Ess88, McB91, MM89]. interface [WBT89, WD96]. International [Ano91b, Ano91a]. interpolating [MC96]. interpolation [CM93, MS92]. interpretation [SH95]. intersections [CBA90]. Investigation [HS90, SNtW94]. iPSC [Ess88, McB91, MM89]. iPSC/1 [Ess88]. iPSC/1-VX [Ess88]. iPSC/860 [McB91]. irregular [Hv90]. ISP [vv88]. Issues [Kha92, Kha91]. issuing [HMN^{+}92]. Iterative [PS97, CK92, MS89, TP93].

Japan [Ano91b, HS96, JB95, Sch97]. JPL [SRS^{+}90]. JPL/CIT [SRS^{+}90].

KAP [LCH^{+}88]. kernel [SH95]. Krylov [CK92]. KSR [BHW^{+}93]. KSR-1 [BHW^{+}93]. KSR1 [Sch93]. Kutta [CPS97].

Laboratory [GSv90]. language

[Sol89b, Sol89c, Tsu91]. LAPACK [DD91]. Large [BPA90, BHWZ96, NK90, Oak90, PS92]. large-scale [NK90]. Law [SS95]. level [HH96, WCMJ89]. libraries [DHW92, PHH91, Van88a]. library [DD91, Du 90]. Light [Hv90]. Linear [DV92, DMR91, DD91, MS89, Ove91b, PZN96, RRS88, SK91, Van90b, Zil93, Zil96]. LINPACK [BDL91]. List [DS96, DS97, Sim95]. load [UT91]. load/store [UT91]. local [MS92, Zha89]. low [HH96]. low-level [HH96]. LU [LC97].

machine [CV93]. machines

[Gar90, KNG^{+}93, Lan88, PZN96, ZWS89, ZWMV92]. magnetically [ACM^{+}91]. mainframes [LCHJ92]. manipulating [OYI91]. Markov [TP93]. Massively [MST96]. mathematical [NH91]. matrices [Hen89, Pin94]. matrix [Lt93, WE91]. Measuring [Haa89]. medicine [Wit91]. medium [PSNK96]. meets [Mir94]. Melbourne [Rui91]. Memory [NMW93, ACK94, CV93, Gen89, Get95, KNG^{+}93, PdR91, PZN96, Zha89]. message [LC97, WD96]. method [CS90, CV93, FZM91, Lt93, MS90a, SSL88, TM90, VD92a, Zil93, Zil96]. methods [CK92, MS90b, TP93, Zha89]. microprocessor [Sv96]. migrating [GKMR89]. migration [HG89]. MIMD [KNG^{+}93, Sol89b, Sol89c]. MIMD/SIMD [Sol89b, Sol89c]. mini [Hv90]. mini- [Hv90]. minisupercomputers [Gen88]. MIPS [Sv96]. model [LJ88, PSNK96, RD92, de 89].

Modeling [REKP94, Coo90, IG91]. Modelling [KC95, JH90]. models [BHWZ96, ZWMV92]. Molecular [BDB93, KNG^{+}93]. molecules [WM91]. MP [SNtW94, RNHW90, WTW90a, WTW90b]. MP/432 [MS90a]. MPI [NAW^{+}96, WD96]. MPQS [LtW88]. MPQS-factoring [LtW88]. multi [IG91]. multi-transputer [IG91]. multicomputer [Rui91]. Multigrid [Hem84, Mic93, MS90b, Sch93]. multiple [BMM97]. multiplying [Hen89]. multiprocessor [ACK94, CM88, RRS88, SH97]. multiprocessors [PdR91, Zha89]. multitasked [WTW90a, WTW90b]. multithreaded [HMN^{+}92]. music [Mir94]. MVS [LHHJ91]. Myrias [BW90].

NAG [Du 90, DHW92]. NAS [SS95]. natural [TM90]. nCube [Pin94]. NEC [EM91, LtW88, MSW92b]. nested [Pin92]. network [ALM96, Kha92, MC96, NK90, Zil96]. newly [Van96, Van97]. Non [MB90, BMM97, MC96]. non-gridded [MC96]. non-standard [BMM97]. Non-traditional [MB90]. nonlinear [NK90, Zil96]. nonsymmetric [Zil96]. note [DHRW92]. November [Ano91b]. NT [SH97]. number [IK90, Kle89, WGOY91]. numbers [Pin94, van90a]. numerical [DD91, GW91]. Nyström [CPS97].

O [LHHJ91, VD92a]. Occam [Hul89]. ODIN [SG90]. off [MS97, Van92a]. Omega [HG89]. Omega-X [HG89]. onto [GKMR89]. operating [LHHJ91]. operational [Van92a]. operations [DMR91]. optimisation [DHNP95]. Optimization [LtW88, RNHW90, Coi90, PL90, RTY90]. Optimizing [BP97, ZWMV92]. optoelectronic [Rui91]. order [HS90, Ove91b]. ordinates [DHRW92]. output [LCH^{+}88]. overview [Wit91].

Parallel [BMM97, CM93, Gen89, KNG^{+}93, Mic93, MST96, MS89, MS90b, MS92, NK90, OYWK91, PdR91, PL90, TP93, TVV96, Zha89, Zil93, Zil96, ACK94, BW90, BHW^{+}93, BDL91, CBA90, CK92, CPS97, DNT96, DHNP95, GH89, KC95, LC97, McB91, MEGS97, Mod89, MC96, O’N88, Oak90, Ove91a, Pan90, Pin92, PS92, SH95, SS95, SDV89, TP89, VD92a, Van90b, ZPN95, ZWMV92, Zoi87, van90a]. Parallelism [Rei87]. Parallelization [RD92, Sch93]. parallelizer [BGT89]. PARAM [PSNK96]. parametric [MS92]. PARFES [Zoi87]. ParkBench [DH95]. ParStone [Was93]. partial [OP93]. Particle [DNT96, Dec88]. Particle-in-cell [DNT96]. particles [Hv90]. Pascal [Tsu91]. Passing [WD96, LC97]. PC [SH97]. peak [BDL91]. Performance [DV92, DS96, DS97, DHNP95, EH95, Get95, Har96, Har97, LCHJ92, LCH^{+}96, LC97, MEGS97, MSW92a, MSW92b, NH91, NMW93, PHH91, Sim95, Sv96, SDMS97, WH90, van91, AHJS90, BP93, Ben89, BDL91, BHWZ96, DD91, EM91, Hoc92, LJ88, LHHJ91, MDL^{+}89, McB93, MS95, Mir94, Ove92b, Roy92]. performances [BMM97]. pictures [MM89]. pipeline [GS90]. Pipelining [PZN96]. PISO [CS90]. plasma [Dec88]. plasmas [ACM^{+}91]. pollution [BHWZ96, ZWMV92]. Polytechnic [ALM96]. Portable [van90a, DD91]. power [Ove91a, NMW93]. POWER/4 [NMW93]. prediction [MEGS97, RTY90, Van92a]. Preliminary [van92b, BP93]. preprocessors [LCH^{+}88]. preserving [CM93]. principles [REKP94]. problem [CM88]. problems [CS90, Lt93, NK90, SK91, TP93, WGOY91]. Proceedings [Ano91b]. process [AKB^{+}97]. processes [Pud89]. processing [RRS88, SRS^{+}90]. processor [ALM96, HMN^{+}92, IK90, LJ88, O’N88, PHH91]. processors [Dow92, Oak90, Ove92a, WCMJ89, van91]. program [Haa89, RNHW90]. programming [ACK94, TP89]. programs [GH89, GKMR89, KC95, Pan90, van92b]. project [ALM96, JH90, Rui91]. projection [Zil93]. Pseudorandom [Kle89]. PVM [Mic93, van93a].

QR [LC97]. quantum [NMW93].

R8000 [Sv96]. radix [Dow92, van91]. radix-2 [van91]. radix-3 [van91]. Random [IK90, van90a]. range [PSNK96]. rapid [Ove91a]. Rayleigh [TVV96]. Rayleigh-Benard [TVV96]. reconstruction [RNHW90]. recurrence [HS90, Ove91b, SK91]. Red [van93a]. relation [Ove91b]. remarks [Hen89]. research [AKB^{+}97]. Resources [NAW^{+}96]. results [Gen89, Van93b, van92b, vd92b]. reuse [SNtW94]. revisited [Ove92b]. RF [WTW90a, WTW90b]. ring [BDB93]. RISC [DMR91]. row [Zil93]. row-projection [Zil93]. run [Hem84]. Runge [CPS97]. Running [ZWS89, BHWZ96].

S [SNtW94]. S-MP [SNtW94]. S3800 [Van93b]. scale [NK90]. Scaling [McB93]. scattering [Hv90]. scheduling [CM88]. Schrödinger [GS90]. SCI [SH97]. science [Got91]. sciences [NK90, Pan91]. scientific [Gar90, KC95, LCHJ92, LCH^{+}96, OYWK91, PHH91, Roy92]. Sea [Coo90]. second [Gen88]. semi [MS89]. semi-asynchronous [MS89]. semiconductor [Lt93]. SERC [GSv90]. series [Ove91a]. shallow [McB93, de 89]. shape [CM93]. shape-preserving [CM93]. Shared [NMW93, ACK94, Gen89, OYI91, Zha89]. shared-memory [ACK94]. shooting [BMM97]. Short [MH92, Van96, Van95a, Van97]. Short-distance [MH92]. Siemens [DHW92, PHH91]. SIMD [BHWZ96, Sol89b, Sol89c]. similarity [Hoc95, MEGS97]. Simple [Ess88, Lan88, SH95]. Simulation [Got92, BDB93, BPA90, CLHP93, Dec88, DNT96, KC95, Pud89, WE91]. simulations [KNG^{+}93, WTW90a, WTW90b]. simultaneous [HMN^{+}92]. Sites [DMS95, DMS97, SB96, Sim97, DMS96]. Six [WGOY91]. social [NK90, Pan91]. Software [DV92, GK95, GW91, Paj90, SG90, Sol89a, Van88a, Van88b]. solidification [CLHP93]. Solution [TM90, MS89, PS97]. solve [TP93]. solver [PZN96, Sch93, WE91]. Solving [GS90, Lt93, SK91, CS90, Zil93]. Some [Hen89, Lt93, PHH91, van92b]. SOR [Zha89]. sort [Dow92]. Sorting [Dow92, Ove90, Ove92a]. SP [LC97]. SP-2 [LC97]. spaced [ACK94]. sparse [PS92, Pin94, Van90b, WE91, Zil93, Zil96]. sparse-matrix [WE91]. spectral [PSNK96, RD92]. SPICE [PdR91]. SPS [BW90]. SPS-2 [BW90]. STAB [FZM91]. stability [ACM^{+}91]. Standard [DV92, BMM97, Fis90, WD96]. statistical [SS95]. Status [Van95b, van93c]. store [UT91]. strategy [MDL^{+}89]. structured [Tsu91]. studies [REKP94]. study [EM91, GKMR89, LCH^{+}88, Nob89, RTY90]. subdivision [CBA90]. subroutines [PHH91]. subspace [CK92]. Summary [vv88, vd92b]. SUPERB [BGT89]. Supercomputer [CLHP93, DMS95, DMS97, KSS96, SG90, SB96, Sim97, DMS96, FGHv89, MB88, Pud89, SH95]. Supercomputers [GK95, HB90, MST96, Ric91, Wit91, GW91, Hv90, Hem84, Kle89, PS97, Ruh90, Sch97, SK91]. Supercomputing [ACH^{+}89, Ano91a, Ano91b, Bar88, Buz90, Du 90, Hul89, JB95, Nie90, Pan91, Wal90, ASN^{+}91, Coi90, GSv90, Kha91, Kha92, Van92a]. superperformance [MB90]. support [Emm89, Pan91]. Supporting [ACK94]. Suprenum [Gil89, TP89, McB93, AHJS90, BGT89, KOS^{+}91, McB91, Sol89a, Sol89b, Sol89c, Tro89, WBT89]. SUPRENUM-1 [McB93, McB91]. Suprenum-Fortran [Sol89b, Sol89c]. Surface [CBA90]. surfaces [MS92]. Swell [Van92a]. SX [EM91, LtW88, LJ88, MSW92b]. SX-2 [LtW88, LJ88]. SX-3 [EM91, MSW92b]. SX-3/44 [MSW92b]. symbolic [Mur91, Paj90]. Symposium [Ano91a, Ano91b]. synchronous [TP93]. synthetic [OP93]. System [SNtW94, Emm89, Gil89, Got92, LHHJ91, Zoi87, DMR91, NMW93]. system-independent [Emm89]. System/6000 [DMR91]. Systems [MS97, Gen89, MEGS97, MS89, PS97, PHH91, Van90b, ZPN95, Zil93, Zil96].

T3D [PS97]. T3E [PS97]. T9000 [BP93]. T9000-a [BP93]. takes [MS97]. task [Hv90]. TERPSICHORE [ACM^{+}91]. tests [vv88]. Three [SRS^{+}90, ZWS89]. Three-dimensional [SRS^{+}90]. time [TVV96]. timings [Dec88]. TOP25 [SB96, Sim97]. TOP500 [DMS95, DS96, DMS96, DS97, DMS97, Sim95, Van96, Van95a, Van97]. traditional [MB90]. Transforms [WH90]. transient [CS90]. transport [DHRW92, Pud89]. transputer [Got92, Hul89, IG91, Zil96]. transputers [BDL91, GS90, Wal90]. triangular [Hen89]. tridiagonal [PS97]. turbulence [BPA90]. two [DHW92, Hen89, LCH^{+}88].

U.S. [DS96, DS97]. UK [AKB^{+}97, JH90]. Unisys [vv88]. universitie [JH90]. University [BW90, Rui91]. Updating [Mod89]. Usage [MS97]. User [VD92a, BW90, WBT89]. users [Emm89].

Using [DV92, Rei87, CBA90, Coo90,

DHRW92, Lt93, Mic93, MSW92a, MSW92b].

V2.0 [vd92b]. V3.0 [van92b]. VAMPIR [NAW^{+}96]. Variations [HH96]. Various [DV92, Zha89]. VAST [LCH^{+}88]. VAST-2 [LCH^{+}88]. Vector [PS92, Dow92, Emm89, IK90, LJ88, LCH^{+}88, OYI91, Ove92a, PHH91, RRS88, UT91, Van90b, WCMJ89, ZWS89, van91]. vectorcomputers [WH90]. Vectorial [Van88a]. vectorizable [Ove90]. Vectorization [CS90, Mur91, SSL88, Van88b, Haa89, Ove91b]. Vectorized [WTW90a, WTW90b]. vectorizing [Nob89, Tsu91]. versus [LC97, Van89]. VF [GKMR89, HG89]. via [Coi90, CM93, Paj90]. Visualization [FZM91, NAW^{+}96, OYWK91]. Visualizing [Pan90]. VLIW [Fis90]. VP [DHW92]. VPP500 [DNT96]. VX [Ess88].

water [McB93, de 89]. weather [PSNK96]. Windows [SH97]. workload [MEGS97]. workstation [ASN^{+}91, LCH^{+}96]. workstations [LCHJ92].

X [HG89, WTW90a, WTW90b]. X-MP [WTW90a, WTW90b].

Y-MP [MS90a, RNHW90]. Y-MP/432 [MS90a]. year [MS95, Pan91]. yields [SH95].

References

Anderson:1989:SD

[ACH^{+}89] B. S. Anderson, D. J. Christensen, N. C. W. Hansen, T. Jordt, J. Moth, and J. Wasniewski. Supercomputing in Denmark. Supercomputer, 6(1): 8–12, January 1989. CODEN SP-COEL. ISSN 0168-7875.

Alaghband:1994:SSM

[ACK94] G. Alaghband, B. Catalucci, and R. Kalathil. Supporting sharedmemory parallel programming environment on a global address spaced multiprocessor. Supercomputer, 11 (1):24–40, April 1994. CODEN SP-COEL. ISSN 0168-7875. Issue labeled April 1994 – April 1995.

Anderson:1991:TCS

[ACM^{+}91] D. V. Anderson, W. A. Cooper, S. Merazzi, U. Schwenn, and R. Gruber. The TERPSICHORE code for the stability analysis of magnetically confined fusion plasmas. Supercomputer, 8(3):32–35, May 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Ashauer:1990:SFC

[AHJS90] R. Ashauer, T. Hoppe, G. Jost, and K. Solchenbach. The Suprenum Fortran compiler — architecture and performance. Supercomputer, 7(3): 19–25, May 1990. CODEN SP-COEL. ISSN 0168-7875.

Allan:1997:HUA

[AKB^{+}97] R. J. Allan, K. Kleese, I. J. Bush, A. Sunderland, Guest, and M. F. HPCI in the UK: from academic research to industrial process. Supercomputer, 13(3–4):4–22, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Aktouf:1996:APN

[ALM96] C. Aktouf, Y. Latrous, and G. Mazare. Asynchronous processor network— the project of the Polytechnic Institute of Grenoble. Supercomputer, 12(3):21–26, August 1996. CODEN SPCOEL. ISSN 0168-7875.

Anonymous:1991:ISS

[Ano91a] Anonymous. International Symposium on Supercomputing ’91. Supercomputer, 8(6):??, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Anonymous:1991:PIS

[Ano91b] Anonymous, editor. Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukuoka, Japan, November 1991. ISBN 4-87378-284-8. LCCN QA76.88.I1991. Also published in/as Supercomputer, volume 8, number 6 (1991).

Apduhan:1991:EAD

[ASN^{+}91] B. O. Apduhan, T. Sueyoshi, Y. Namiuchi, T. Tezuka, and I. Arita. Experiments and analysis of distributed supercomputing in a distributed workstation environment. Supercomputer, 8(6):90–100, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Barry:1988:SUA

[Bar88] J. Barry. Supercomputing down under (australian facilities). Supercomputer, 5(5):9–14, September 1988. CODEN SPCOEL. ISSN 0168-7875.

Bekker:1993:MDS

[BDB93] H. Bekker, E. J. Dijkstra, and H. J. C. Berendsen. Molecular dynamics simulation on an i860 based ring architecture. Supercomputer, 10 (2):4–10, March 1993. CODEN SP-COEL. ISSN 0168-7875.

Bisseling:1991:TPP

[BDL91] R. H. Bisseling, L. Daniel, and J. C. Loyens. Towards peak parallel LIN-PACK performance on 400 transputers. Supercomputer, 8(5):20–27, September 1991. CODEN SPCOEL. ISSN 0168-7875.

Bennett:1989:CEC

[Ben89] D. Bennett. CERFACS: a European center for high performance computation. Supercomputer, 6(1):13–16, January 1989. CODEN SPCOEL. ISSN 0168-7875.

Bast:1989:SSP

[BGT89] H.-J. Bast, M. Gerndt, and C.-A.

Thole. SUPERB — the Suprenum parallelizer. Supercomputer, 6(2): 51–57, March 1989. CODEN SP-COEL. ISSN 0168-7875.

Bendtsen:1993:EKP

[BHW^{+}93] C. Bendtsen, P. C. Hansen, J. Wasniewski, J. B. Hansen, and J. N. Sorensen. Experience with the KSR-1 parallel computer. Supercomputer, 10(6):34–43, ???? 1993. CODEN SPCOEL. ISSN 0168-7875.

Brown:1996:CPS

[BHWZ96] J. Brown, P. C. Hansen, J. Wasniewski, and Z. Zlatev. Comparing the performance of SIMD computers by running large air pollution models. Supercomputer, 12(2):21– 35, March 1996. CODEN SPCOEL. ISSN 0168-7875.

Bellavia:1997:PPN

[BMM97] S. Bellavia, M. Macconi, and B. Morini. Parallel performances of a non-standard multiple shooting. Supercomputer, 13(3–4):45–56, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Bader:1993:TPE

[BP93] G. Bader and B. Przywara. T9000-a preliminary evaluation of arithmetic performance. Supercomputer, 10(2): 26–34, March 1993. CODEN SP-COEL. ISSN 0168-7875.

Benkner:1997:OHA

[BP97] S. Benkner and M. Pantano. Optimizing HPF for advanced applications. Supercomputer, 13(2):31–43, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Benocci:1990:LES

[BPA90] C. Benocci, A. Pinelli, and A. Abba. Large eddy simulation of turbulence on an Alliant FX computer. Supercomputer, 7(6):77–87, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Buzbee:1990:SE

[Buz90] B. Buzbee. Supercomputing and the environment. Supercomputer, 7(2): 7–17, March 1990. CODEN SP-COEL. ISSN 0168-7875.

Baker:1990:FUE

[BW90] D. Baker and T. Wilson. First user experiences on the Myrias SPS-2 parallel computer at the University of Calgary. Supercomputer, 7(5):6– 11, September 1990. CODEN SP-COEL. ISSN 0168-7875.

Chang:1990:SIU

[CBA90] Long Chyr Chang, W. Bein, and E. Angel. Surface intersections using parallel subdivision. Supercomputer, 7(6):62–70, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Chronopoulos:1992:TEP

[CK92] A. Chronopoulos and S. K. Kim.

Towards efficient parallel implementation of Krylov subspace iterative methods. Supercomputer, 9(1):4–17, January 1992. CODEN SPCOEL. ISSN 0168-7875.

Crempien-Laborie:1993:SSS

[CLHP93] J. E. Crempien-Laborie, J. C.

Heinrich, and D. Poirier. Supercomputer simulation of the solidification of binary alloys. Supercomputer, 10 (1):15–23, January 1993. CODEN SPCOEL. ISSN 0168-7875.

Cohen:1988:HAM

[CM88] E. Cohen and R. Miller. Hypercube algorithms for the multiprocessor scheduling problem. Supercomputer, 5(5):17–32, September 1988. CODEN SPCOEL. ISSN 0168-7875.

Conti:1993:PSP

[CM93] C. Conti and R. Morandi. Parallel shape-preserving interpolation via Bézier curves. Supercomputer, 10 (6):24–29, ???? 1993. CODEN SP-COEL. ISSN 0168-7875.

Coiffier:1990:IDA

[Coi90] B. Coiffier. Integrated design, analysis and optimization via supercomputing. Supercomputer, 7(2):55–64, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Cooper:1990:SMU

[Coo90] A. Cooper. Sea modeling using an AMT DAP. Supercomputer, 7(2): 65–71, March 1990. CODEN SP-COEL. ISSN 0168-7875.

Crisci:1997:PRK

[CPS97] M. R. Crisci, B. Paternoster, and G. Sigillo. A parallel Runge–Kutta– Nyström code for y′′ = f(t, y). Supercomputer, 13(3–4):57–66, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Chuan:1990:VPM

[CS90] C. H. Chuan and W. Schreiber. Vectorization of the PISO method for solving 3-D, transient heat convection problems. Supercomputer, 7(5): 45–51, September 1990. CODEN SPCOEL. ISSN 0168-7875.

Crone:1993:CAC

[CV93] L. G. C. Crone and H. A. Van der Vorst. Communication aspects of the conjugate gradient method on a distributed-memory machine. Supercomputer, 10(6):4–9, ???? 1993. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1991:LPH

[DD91] J. Dongarra and J. Demmel. LA-PACK: a portable high-performance numerical library for linear algebra. Supercomputer, 8(6):33–38, November 1991. CODEN SPCOEL. ISSN 0168-7875.

deGoede:1989:CMS

[de 89] E. de Goede. A computational model for 3-D shallow water flows on the Alliant FX/4. Supercomputer, 6 (4):43–49, July 1989. CODEN SP-COEL. ISSN 0168-7875.

Decyk:1988:BTP

[Dec88] V. Decyk. Benchmark timings with particle plasma simulation codes.

Supercomputer, 5(5):33–43, September 1988. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1995:PBC

[DH95] J. J. Dongarra and T. Hey.

The ParkBench benchmark collection. Supercomputer, 11(2–3):94– 114, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Dunlop:1995:PEP

[DHNP95] A. N. Dunlop, A. J. G. Hey, D. A.

Nicole, and D. J. Pritchard. Performance estimating for parallel performance optimisation. Supercomputer, 11(4):19–30, September 1995. CO-DEN SPCOEL. ISSN 0168-7875.

Datta:1992:NDE

[DHRW92] R. P. Datta, A. S. Hira, A. K. Ray, and B. R. Wienke. A note on 2-D electron transport using discrete ordinates. Supercomputer, 9(4):15–21, July 1992. CODEN SPCOEL. ISSN 0168-7875.

DuCroz:1992:CTF

[DHW92] J. Du Croz, J. S. Hesthaven, and J. Wasniewski. Comparison of two FFT libraries on the Amdahl/ Fujitsu VP computer — NAG and Siemens libraries. Supercomputer, 9 (5):31–37, September 1992. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1991:IRS

[DMR91] J. J. Dongarra, P. Mayes, and G. Radicati di Brozolo. The IBM RISC System/6000 and linear algebra operations. Supercomputer, 8 (4):15–30, July 1991. CODEN SP-COEL. ISSN 0168-7875.

Dongarra:1995:TSS

[DMS95] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 11(2–3): 133–163 (or 164–194??), June 1995. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1996:TSS

[DMS96] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 12(1): 91–120, January 1996. CODEN SP-COEL. ISSN 0168-7875.

Dongarra:1997:TSS

[DMS97] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 13(1): 89–120, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Decyk:1996:PSF

[DNT96] V. K. Decyk, H. Naitou, and S. Tokuda. Particle-in-cell simulation on the Fujitsu VPP500 parallel computer. Supercomputer, 12(4): 28–32, December 1996. CODEN SP-COEL. ISSN 0168-7875.

Dow:1992:SVP

[Dow92] M. Dow. Sorting on vector processors — the binary radix sort. Supercomputer, 9(1):18–26, January 1992. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1996:HPC

[DS96] J. J. Dongarra and H. D. Simon. High performance computing in the U.S. in 1995 — an analysis on the basis of the TOP500 list. Supercomputer, 12(1):16–22, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1997:HPC

[DS97] J. J. Dongarra and H. D. Simon.

High performance computing in the U.S. in 1996 — an analysis on the basis of the TOP500 list. Supercomputer, 13(1):19–28, ???? 1997. CO-DEN SPCOEL. ISSN 0168-7875.

DuCroz:1990:SNL

[Du 90] J. Du Croz. Supercomputing with the NAG library. Supercomputer, 7 (2):72–80, March 1990. CODEN SP-COEL. ISSN 0168-7875.

Dongarra:1992:PVC

[DV92] J. J. Dongarra and H. A. Van der Vorst. Performance of various computers using standard linear equations software in a Fortran environment. Supercomputer, 9(5):17–30, September 1992. CODEN SPCOEL. ISSN 0168-7875.

Deconinck:1997:IEH

[DVB^{+}97] G. Deconinck, T. Varvarigou, O. Botti, D. de Florio, A. Kontizas, M. Truyens, W. Rosseel, R. Lauwereins, F. Cassinari, S. Graeber, and U. Knaak. Industrial embedded HPC applications. Supercomputer, 13(3–4):23–44, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Emmen:1995:HPC

[EH95] A. Emmen and U. Harms. High performance computing in Europe. Supercomputer, 11(2–3):46–60, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Eoyang:1991:NSP

[EM91] C. Eoyang and R. Mendez. NEC SX-3 performance study. Supercomputer, 8(4):31–37, July 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Emmen:1989:TSI

[Emm89] A. Emmen. Towards a systemindependent approach to the support of vector computer users. Supercomputer, 6(6):41–46, November 1989. CODEN SPCOEL. ISSN 0168-7875.

Ess:1988:SCI

[Ess88] M. Ess. Simple convolution on the Intel iPSC/1 and iPSC/1-VX. Supercomputer, 5(4):22–30, July 1988. CODEN SPCOEL. ISSN 0168-7875.

Friedli:1989:ESB

[FGHv89] A. Friedli, W. Gentzsch, R. Hockney, and A. van der Steen. A European supercomputer benchmark effort. Supercomputer, 6(6):14–17, November 1989. CODEN SPCOEL. ISSN 0168-7875.

Fisher:1990:VAI

[Fis90] J. Fisher. VLIW architectures: an inevitable standard for the future? Supercomputer, 7(2):29–36, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Fujino:1991:VCB

[FZM91] S. Fujino, Shaoliang Zhang, and M. Mori. Visualization of convergence behavior of the Bi-CG STAB method. Supercomputer, 8(6):127– 135, November 1991. CODEN SP-COEL. ISSN 0168-7875.

Garrett:1990:AMS

[Gar90] P. Garrett. Abstract machines for scientific computation. Supercomputer, 7(2):37–44, March 1990. CO-DEN SPCOEL. ISSN 0168-7875.

Gentzsch:1988:SGM

[Gen88] W. Gentzsch. The second generation of minisupercomputers. Supercomputer, 5(4):16–21, July 1988. CO-DEN SPCOEL. ISSN 0168-7875.

Gentzsch:1989:PBR

[Gen89] W. Gentzsch. Parallel benchmark results for shared memory systems. Supercomputer, 6(4):10–16, July 1989. CODEN SPCOEL. ISSN 0168-7875.

Getov:1995:PCC

[Get95] V. S. Getov. Performance characterisation of the cache memory effect. Supercomputer, 11(4):31–49, September 1995. CODEN SPCOEL. ISSN 0168-7875.

Griffin:1989:EED

[GH89] James Griffin and Robert Hiromoto.

Experience with an experimental debugger for parallel programs. Supercomputer, 6(5):44–50, September 1989. CODEN SPCOEL. ISSN 0168-7875.

Giloi:1989:SS

[Gil89] W. Giloi. Suprenum — the system.

Supercomputer, 6(2):13–19, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Geiger:1995:ASS

[GK95] A. Geiger and N. Kroll. Application software for supercomputers. Supercomputer, 11(2–3):115–132, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Guest:1989:SMC

[GKMR89] M. Guest, J. Kendrick, R. Mathie, and B. Ralston. A study into migrating computational chemistry programs onto an IBM 3090/VF. Supercomputer, 6(5):12–16, September 1989. CODEN SPCOEL. ISSN 0168-7875.

Goto:1991:CSE

[Got91] E. Goto. Computational science and engineering. Supercomputer, 8(6):5– 11, November 1991. CODEN SP-COEL. ISSN 0168-7875.

Gottwald:1992:SAD

[Got92] S. Gottwald. Simulation of atmospheric dispersion on a transputer system. Supercomputer, 9(1):43–50, January 1992. CODEN SPCOEL. ISSN 0168-7875.

Gallagher:1990:SIS

[GS90] A. Gallagher and S. Scott. Solving the inhomogeneous Schrödinger equation on a bi-directional pipeline of transputers. Supercomputer, 7(6): 88–97, November 1990. CODEN SP-COEL. ISSN 0168-7875.

Guest:1990:CSS

[GSv90] M. Guest, P. Sherwood, and J. van Lenthe. Concurrent supercomputing at SERC Daresbury Laboratory. Supercomputer, 7(2):89–103, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Geers:1991:HEB

[GW91] N. Geers and W. Walde. Highly efficient basic numerical software for supercomputers. Supercomputer, 8 (6):136–145, November 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Haas:1989:MDV

[Haa89] U. Haas. Measuring the degree of vectorization by an artificial benchmark program. Supercomputer, 6 (4):26–36, July 1989. CODEN SP-COEL. ISSN 0168-7875.

Harms:1996:HPC

[Har96] U. Harms. High performance computing in Europe. Supercomputer, 12(1):30–37, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Harms:1997:HPC

[Har97] U. Harms. High performance computing in Europe. Supercomputer, 13(1):38–44, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Howard:1990:SG

[HB90] S. Howard and G. Byrd. Supercomputers and galaxies. Supercomputer, 7(3):36–49, May 1990. CODEN SP-COEL. ISSN 0168-7875.

Hemker:1984:MAR

[Hem84] P. W. Hemker. Multigrid algorithms run on supercomputers. Supercomputer, 4(????):44–51, 1984. CODEN SPCOEL. ISSN 0168-7875.

Heng:1989:SRM

[Hen89] A.-K. Heng. Some remarks on multiplying two triangular matrices. Supercomputer, 6(5):28–33, September 1989. CODEN SPCOEL. ISSN 0168-7875.

Hague:1989:OMI

[HG89] J. Hague and F. Goloway. 3-D Omega-X migration on the IBM 3090/VF. Supercomputer, 6(5):17– 27, September 1989. CODEN SP-COEL. ISSN 0168-7875.

Hernandez:1996:VLL

[HH96] E. Hernandez and T. Hey. Variations on low-level communication benchmarks. Supercomputer, 12(4): 16–27, December 1996. CODEN SP-COEL. ISSN 0168-7875.

Hirata:1992:MPA

[HMN^{+}92] H. Hirata, Y. Mochizuki, A. Nishmura, Y. Nakase, and T. Nishizawa. A multithreaded processor architecture with simultaneous instruction issuing. Supercomputer, 9(3):23–39, May 1992. CODEN SPCOEL. ISSN 0168-7875.

Hockney:1992:FBP

[Hoc92] R. Hockney. A framework for benchmark performance analysis. Supercomputer, 9(2):9–22, March 1992. CODEN SPCOEL. ISSN 0168-7875.

Hockney:1995:CS

[Hoc95] R. W. Hockney. Computational similarity. Supercomputer, 11(4):102– 123, September 1995. CODEN SP-COEL. ISSN 0168-7875.

Hafner:1990:IDA

[HS90] H. Hafner and W. Schonauer. Investigation of different algorithms for the first order recurrence. Supercomputer, 7(6):34–41, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Hoffmann:1996:DJ

[HS96] G.-R. Hoffmann and E. Schnepf. Developments in Japan. Supercomputer, 12(1):23–29, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Hull:1989:STO

[Hul89] E. Hull. Supercomputing with the transputer and Occam. Supercomputer, 6(6):18–23, November 1989. CODEN SPCOEL. ISSN 0168-7875.

Hage:1990:LSI

[Hv90] J. Hage and J. van Kats. Light scattering by inhomogeneous, irregular particles — a task for (mini- )supercomputers. Supercomputer, 7 (4):30–38, July 1990. CODEN SP-COEL. ISSN 0168-7875.

Ierotheou:1991:MTA

[IG91] C. Ierotheou and E. Galea. A multitransputer approach to fire field modeling. Supercomputer, 8(2):16– 24, March 1991. CODEN SPCOEL. ISSN 0168-7875.

Ito:1990:RNG

[IK90] N. Ito and Y. Kanada. Random number generation on a vector processor. Supercomputer, 7(1):29–35, January 1990. CODEN SPCOEL. ISSN 0168-7875.

Jarp:1995:SJ

[JB95] S. Jarp and W. Bez. Supercomputing in Japan. Supercomputer, 11(2– 3):31–45, June 1995. CODEN SP-COEL. ISSN 0168-7875.

James:1990:UUG

[JH90] I. James and B. Hoskins. The UK universities’ global atmospheric modelling project. Supercomputer, 7(2):104–113, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Kempf:1995:MSS

[KC95] G. Kempf and C. Caremoli. Modelling and simulation of scientific programs on parallel architectures. Supercomputer, 11(4):50–62, September 1995. CODEN SPCOEL. ISSN 0168-7875.

Khan:1991:CIS

[Kha91] S. Khan. Current issues in supercomputing. Supercomputer, 8(1): 18–29, January 1991. CODEN SP-COEL. ISSN 0168-7875.

Khan:1992:INS

[Kha92] S. Khan. Issues in network supercomputing. Supercomputer, 9(5):9– 16, September 1992. CODEN SP-COEL. ISSN 0168-7875.

Kleijnen:1989:PNG

[Kle89] J. Kleijnen. Pseudorandom number generation on supercomputers. Supercomputer, 6(6):34–40, November 1989. CODEN SPCOEL. ISSN 0168-7875.

Kalia:1993:PAM

[KNG^{+}93] R. K. Kalia, A. Nakano, D. L. Greenwell, P. Vashishta, and S. W. de Leeuw. Parallel algorithms for molecular dynamics simulations on distributed-memory MIMD machines. Supercomputer, 10(2):11– 25, March 1993. CODEN SPCOEL. ISSN 0168-7875.

Kehl:1991:ABS

[KOS^{+}91] E. Kehl, K.-D. Oertel, K. Solchenbach, R. Vogelsang, and O. McBryan. Application benchmarks on Suprenum. Supercomputer, 8(2):6–15, March 1991. CODEN SPCOEL. ISSN 0168-7875.

Koski:1996:SB

[KSS96] K. Koski, S. Saarinen, and O. Serimaa. Supercomputer benchmarking. Supercomputer, 12(2):48–71, March 1996. CODEN SPCOEL. ISSN 0168-7875.

Lang:1988:SCF

[Lan88] U. Lang. Simple CPU-benchmarks in Fortran 77 on different machines. Supercomputer, 5(6):26–33, November 1988. CODEN SPCOEL. ISSN 0168-7875.

Luecke:1997:HPF

[LC97] G. R. Luecke and J. J. Coyle. High Performance Fortran versus explicit message passing on the IBM SP-2 for the parallel LU, QR, and Cholesky factorizations. Supercomputer, 13(2):4–14, ???? 1997. CO-DEN SPCOEL. ISSN 0168-7875.

Luecke:1988:CSK

[LCH^{+}88] G. Luecke, J. Coyle, W. Haque, J. Hoekstra, H. Jespersen, and R. Schmidt. A comparative study of KAP and VAST-2: two automatic vector preprocessors with Fortran 8x output. Supercomputer, 5(6):15–25, November 1988. CODEN SPCOEL. ISSN 0168-7875.

Luecke:1996:PCW

[LCH^{+}96] G. Luecke, J. Coyle, W. Haque, J. Hoekstra, and H. Jespersen. Performance comparison of workstation clusters for scientific computing. Supercomputer, 12(2):4–20, March 1996. CODEN SPCOEL. ISSN 0168-7875.

Luecke:1992:PWM

[LCHJ92] G. Luecke, J. Coyle, J. Hoekstra, and H. Jesperson. Performance of workstations and mainframes for scientific computing. Supercomputer, 9(6):10–19, November 1992. CODEN SPCOEL. ISSN 0168-7875.

Luecke:1991:CPE

[LHHJ91] G. Luecke, W. Haque, J. Hoekstra, and H. Jespersen. I/O considerations for performance enhancement under the MVS operating system. Supercomputer, 8(5):41–50, September 1991. CODEN SPCOEL. ISSN 0168-7875.

Liu:1988:PES

[LJ88] Yiwei Liu and O. Johnson. A performance evaluation of the SX-2 vector processor based on a new performance model. Supercomputer, 5 (6):34–52, November 1988. CODEN SPCOEL. ISSN 0168-7875.

Lai:1993:SSD

[Lt93] C.-H. Lai and H. J. J. te Riele. Solving some 1-D semiconductor device problems on a matrix coprocessor using a domain decomposition method. Supercomputer, 10(1): 24–32, January 1993. CODEN SP-COEL. ISSN 0168-7875.

Lioen:1988:OMA

[LtW88] W. Lioen, H. te Riele, and D. Winter. Optimization of the MPQSfactoring algorithm on the Cyber 205 and the NEC SX-2. Supercomputer, 5(4):42–50, July 1988. CO-DEN SPCOEL. ISSN 0168-7875.

Messina:1988:CCS

[MB88] P. Messina and C. Baillie. Caltech concurrent supercomputer facilities. Supercomputer, 5(5):7–8, September 1988. CODEN SPCOEL. ISSN 0168-7875.

Mantell:1990:NTA

[MB90] S. Mantell and M. Burwen. Nontraditional applications for superperformance computing. Supercomputer, 7(2):114–119, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Morandi:1996:PAC

[MC96] R. Morandi and C. Conti. A parallel algorithm to construct a curve network interpolating non-gridded functional data in R^{3}. Supercomputer, 12(4):33–45, December 1996. CODEN SPCOEL. ISSN 0168-7875.

McBryan:1991:CII

[McB91] O. McBryan. A comparison of the Intel iPSC/860 and the Suprenum-1 parallel computers. Supercomputer, 8(1):6–17, January 1991. CODEN SPCOEL. ISSN 0168-7875.

McBryan:1993:SPS

[McB93] O. A. McBryan. Scaling performance of the shallow water equations on the SUPRENUM-1. Supercomputer, 10(1):4–14, January 1993. CODEN SPCOEL. ISSN 0168-7875.

Markov:1989:HPC

[MDL^{+}89] S. Markov, P. Daskalov, V. Lazarov, O. Kostadinoc, B. Janakiev, P. Panteleev, S. Palikarev, K. Kirov, and T. Velitchkov. High performance computer complex — strategy and evolution. Supercomputer, 6(3):12– 23, ???? 1989. CODEN SPCOEL. ISSN 0168-7875.

Meajil:1997:PPP

[MEGS97] A. I. Meajil, T. El-Ghazawi, and T. Sterling. Performance prediction of parallel systems based on workload similarity. Supercomputer, 13 (2):15–30, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Moth:1992:SDC

[MH92] J. Moth and P. C. Hansen. Shortdistance communication on the CM-200. Supercomputer, 9(5):38–43, September 1992. CODEN SPCOEL. ISSN 0168-7875.

Michielse:1993:PMU

[Mic93] P. Michielse. Parallel multigrid using PVM. Supercomputer, 10(6):10– 23, ???? 1993. CODEN SPCOEL. ISSN 0168-7875.

Miranda:1994:CCM

[Mir94] E. R. Miranda. Chaosynth — computer music meets high-performance computing. Supercomputer, 11(1): 16–23, April 1994. CODEN SP-COEL. ISSN 0168-7875. Issue labeled April 1994 – April 1995.

Miller:1989:CAD

[MM89] R. Miller and S. Miller. Convexity algorithms for digitized pictures on an Intel iPSC hypercube. Supercomputer, 6(3):45–53, ???? 1989. CO-DEN SPCOEL. ISSN 0168-7875.

Modi:1989:UCF

[Mod89] J. Modi. Updating Choleski factors in parallel. Supercomputer, 6 (1):23–30, January 1989. CODEN SPCOEL. ISSN 0168-7875.

Morandi:1989:PSA

[MS89] R. Morandi and F. Sgallari. Parallel semi-asynchronous algorithms for the iterative solution of linear systems. Supercomputer, 6(4):17–25, July 1989. CODEN SPCOEL. ISSN 0168-7875.

Morandi:1990:CMI

[MS90a] R. Morandi and A. Sestini. Cray Y-MP/432 implementation of a domain decomposition method. Supercomputer, 7(6):27–33, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Morandi:1990:PCM

[MS90b] R. Morandi and A. Sestini. Parallel computing for multigrid methods. Supercomputer, 7(4):39–47, July 1990. CODEN SPCOEL. ISSN 0168-7875.

Morandi:1992:PLB

[MS92] R. Morandi and A. Sestini. Parallel local blending interpolation for parametric surfaces. Supercomputer, 9(1):27–32, January 1992. CODEN SPCOEL. ISSN 0168-7875.

Meuer:1995:YCH

[MS95] H. W. Meuer and E. Strohmaier.

1994 — a year of change (high performance computing). Supercomputer, 11(2–3):6–20, June 1995. CO-DEN SPCOEL. ISSN 0168-7875.

Meuer:1996:NBB

[MS96] H. W. Meuer and E. Strohmaier.

New building blocks for HPC in 1995. Supercomputer, 12(1):6–15, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Meuer:1997:IUH

[MS97] H. W. Meuer and E. Strohmaier.

1996: The industrial usage of HPC systems takes off. Supercomputer, 13 (1):6–18, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Mierendorff:1996:CSA

[MST96] H. Mierendorff, A. Schüller, and U. Trottenberg. Cluster of supercomputers — an alternative to massively parallel computing. Supercomputer, 12(1):81–90, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Moncrieff:1992:PCC

[MSW92a] D. Moncrieff, V. R. Saunders, and S. Wilson. Performance of the C-90 computer in electron correlation calculations using ccMBPT. Supercomputer, 9(3):4–6, May 1992. CODEN SPCOEL. ISSN 0168-7875.

Moncrieff:1992:PNS

[MSW92b] D. Moncrieff, V. R. Saunders, and S. Wilson. Performance of the NEC SX-3/44 computer in electron correlation calculations using ccMBPT. Supercomputer, 9(4):4–7, July 1992. CODEN SPCOEL. ISSN 0168-7875.

Murao:1991:VSD

[Mur91] H. Murao. Vectorization of symbolic determinant calculation. Supercomputer, 8(3):36–48, May 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Nagel:1996:VVA

[NAW^{+}96] W. E. Nagel, A. Arnold, M. Weber, H.-C. Hoppe, and K. Solchenbach. VAMPIR: Visualization and analysis of MPI resources. Supercomputer, 12(1):69–80, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Nagai:1991:PEM

[NH91] T. Nagai and Y. Hatano. Performance evaluation of mathematical functions. Supercomputer, 8(6):46– 56, November 1991. CODEN SP-COEL. ISSN 0168-7875.

Nieminen:1990:SF

[Nie90] R. Nieminen. Supercomputing in Finland. Supercomputer, 7(3):13–18, May 1990. CODEN SPCOEL. ISSN 0168-7875.

Nagurney:1990:PCL

[NK90] A. Nagurney and Dae-Shik Kim.

Parallel computation of large-scale nonlinear network problems in the social and economic sciences. Supercomputer, 7(6):50–61, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Nanayakkara:1993:PCI

[NMW93] A. Nanayakkara, D. Moncrieff, and S. Wilson. Performance characteristics of the IBM Shared Memory System (POWER/4) in a quantum chemical application. Supercomputer, 10(6):30–33, ???? 1993. CODEN SPCOEL. ISSN 0168-7875.

Nobayashi:1989:CSA

[Nob89] H. Nobayashi. A comparative study of automatically vectorizing Fortran compilers. Supercomputer, 6(6):24– 33, November 1989. CODEN SP-COEL. ISSN 0168-7875.

Oakley:1990:CLP

[Oak90] B. Oakley. The challenge of the large parallel processors in Europe. Supercomputer, 7(2):18–28, March 1990. CODEN SPCOEL. ISSN 0168-7875.

ONeill:1988:BGP

[O’N88] G. O’Neill. The Butterfly GP1000 parallel processor and its applications. Supercomputer, 5(6):6–14, November 1988. CODEN SPCOEL. ISSN 0168-7875.

Obe:1993:ISB

[OP93] A. Obe and W. Paul. Individual synthetic benchmarks and partial decompilation. Supercomputer, 10 (4–5):58–71, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

Overill:1990:FFV

[Ove90] R. Overill. A family of fullyvectorizable sorting algorithms. Supercomputer, 7(6):71–76, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Overall:1991:RPE

[Ove91a] R. E. Overall. On the rapid parallel evaluation of power series. Supercomputer, 8(4):38–41, July 1991. CODEN SPCOEL. ISSN 0168-7875.

Overill:1991:EVG

[Ove91b] R. Overill. On the efficient vectorization of the general first-order linear recurrence relation. Supercomputer, 8(2):31–36, March 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Overill:1992:BSV

[Ove92a] R. E. Overill. Bitonic sorting for vector processors. Supercomputer, 9 (2):4–8, March 1992. CODEN SP-COEL. ISSN 0168-7875.

Overill:1992:HPR

[Ove92b] R. E. Overill. Hypercube performance revisited. Supercomputer, 9 (5):4–8, September 1992. CODEN SPCOEL. ISSN 0168-7875.

Ochi:1991:VAM

[OYI91] H. Ochi, S. Yajima, and N. Ishiura. A vector algorithm for manipulating Boolean functions based on shared binary decision diagrams. Supercomputer, 8(6):101–118, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Oyake:1991:PIG

[OYWK91] I. Oyake, T. Yoshida, Y. Wauke, and A. Kawai. Parallel image generation for scientific visualization. Supercomputer, 8(6):75–89, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Pajunen:1990:GME

[Paj90] P. Pajunen. Generation of more efficient Fortran code via symbolic software. Supercomputer, 7(1):6–41, January 1990. CODEN SPCOEL. ISSN 0168-7875.

Pancake:1990:VBP

[Pan90] Cherri M. Pancake. Visualizing the behavior of parallel programs. Supercomputer, 7(5):31–37, September 1990. CODEN SPCOEL. ISSN 0168-7875.

Pancake:1991:SSS

[Pan91] C. Pancake. Supercomputing support for the social sciences: a fiveyear forecast. Supercomputer, 8(3): 14–21, May 1991. CODEN SP-COEL. ISSN 0168-7875.

Pacheco:1991:PSD

[PdR91] P. S. Pacheco, J. M. del Rosario, and T. Rashid. Parallel SPICE on distributed memory multiprocessors. Supercomputer, 8(6):119–126, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Pool:1991:PAS

[PHH91] T. Pool, P. Hasenbusch, and R. Heil. Performance analysis of some subroutines of scientific libraries on Siemens vector processor systems. Supercomputer, 8(1):38–52, January 1991. CODEN SPCOEL. ISSN 0168-7875.

Pini:1992:DDN

[Pin92] G. Pini. Domain decomposition and nested grids in a parallel environment. Supercomputer, 9(4):22–28, July 1992. CODEN SPCOEL. ISSN 0168-7875.

Pini:1994:CCN

[Pin94] G. Pini. Calculation of condition numbers of sparse matrices on the nCube 2. Supercomputer, 11(1):41– 48, April 1994. CODEN SPCOEL. ISSN 0168-7875. Issue labeled April 1994 – April 1995.

Pardalos:1990:PBB

[PL90] P. Pardalos and Xiaoye Li. Parallel branch-and-bound algorithms for combinatorial optimization. Supercomputer, 7(5):23–30, September 1990. CODEN SPCOEL. ISSN 0168-7875.

Pini:1992:VPC

[PS92] G. Pini and F. Sartoretto. Vector and parallel codes for large sparse eigenproblems. Supercomputer, 9 (4):29–39, July 1992. CODEN SP-COEL. ISSN 0168-7875.

Pini:1997:ISB

[PS97] G. Pini and F. Sartoretto. Iterative solution of block tridiagonal systems on the Cray T3D and T3E supercomputers. Supercomputer, 13(3–4): 67–82, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Purohit:1996:GSM

[PSNK96] S. C. Purohit, T. V. Singh, P. S. Narayanan, and A. Kaginalkar. Global spectral medium range weather forecasting model on PARAM. Supercomputer, 12(3):27– 36, August 1996. CODEN SPCOEL. ISSN 0168-7875.

Pudykiewicz:1989:ASS

[Pud89] J. Pudykiewicz. An application of a supercomputer for the simulation of atmospheric transport processes. Supercomputer, 6(3):36–44, ???? 1989. CODEN SPCOEL. ISSN 0168-7875.

Papay:1996:PGL

[PZN96] J. Papay, M. J. Zemerly, and G. R.

Nudd. Pipelining the Givens linear solver on distributed memory machines. Supercomputer, 12(3):37–43, August 1996. CODEN SPCOEL. ISSN 0168-7875.

Rotstayn:1992:PSG

[RD92] L. Rotstayn and M. Dix. Parallelization of a spectral general circulation model. Supercomputer, 9(1):33–42, January 1992. CODEN SPCOEL. ISSN 0168-7875.

Reid:1987:EPU

[Rei87] J. Reid. The exploitation of parallelism by using Fortran 8X features. Supercomputer, 19 (??)(????):8–18, 1987. CODEN SPCOEL. ISSN 0168-7875.

Ross:1994:MBB

[REKP94] R. B. Ross, W. C. Ermler, C. W. Kern, and R. M. Pitzer. Modeling bulk beryllium through first principles cluster studies. Supercomputer, 11(1):49–59, April 1994. CODEN SPCOEL. ISSN 0168-7875. Issue labeled April 1994 – April 1995.

Richards:1991:SDD

[Ric91] G. Richards. Supercomputers in drug discovery. Supercomputer, 8(2): 25–30, March 1991. CODEN SP-COEL. ISSN 0168-7875.

Roberts:1990:ODA

[RNHW90] K. Roberts, A. Niu, Chia Yung Han, and W. Wee. Optimization of a 3-d algebraic reconstruction program on a Cray Y-MP. Supercomputer, 7(5):14–22, September 1990. CODEN SPCOEL. ISSN 0168-7875.

Royer:1992:EPA

[Roy92] D. Royer. Effective performance analysis of scientific algorithms on a hypercube. Supercomputer, 9(4): 8–14, July 1992. CODEN SPCOEL. ISSN 0168-7875.

Radicati:1988:BPL

[RRS88] G. Radicati, Y. Robert, and P. Sguazzero. Block processing in linear algebra on the IBM 3090 vector multiprocessor. Supercomputer, 5(1):15–25, January 1988. CODEN SPCOEL. ISSN 0168-7875.

Ralston:1990:FPS

[RTY90] B. Ralston, F. Thomas, and F. Yeung. Flood prediction — a study in Fortran optimization and connectivity. Supercomputer, 7(4):48–51, July 1990. CODEN SPCOEL. ISSN 0168-7875.

Ruhle:1990:FAS

[Ruh90] R. Ruhle. Fast access to supercomputers. Supercomputer, 7(2):45– 54, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Ruighaver:1991:MUO

[Rui91] A. B. Ruighaver. The Melbourne University optoelectronic multicomputer project. Supercomputer, 8(6): 22–32, November 1991. CODEN SP-COEL. ISSN 0168-7875.

Simon:1996:TSS

[SB96] H. D. Simon and W. Bez. The TOP25 supercomputer sites. Supercomputer, 12(1):38–47, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Schwardmann:1993:PMS

[Sch93] U. Schwardmann. Parallelization of a multigrid solver on the KSR1. Supercomputer, 10(3):4–12, May 1993. CODEN SPCOEL. ISSN 0168-7875.

Schnepf:1997:DJS

[Sch97] E. Schnepf. Developments in Japan [supercomputers]. Supercomputer, 13(1):29–37, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Strohmaier:1997:HPC

[SDMS97] E. Strohmaier, J. J. Dongarra, H. W. Meuer, and H. D. Simon. High-performance computing in industry. Supercomputer, 13(1):74–88, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Sinha:1989:FCF

[SDV89] U. Sinha, M. Deshpande, and S. Varagapalli. Flow computation on the Flosolver parallel computer. Supercomputer, 6(4):37–42, July 1989. CODEN SPCOEL. ISSN 0168-7875.

Schreiner:1990:SSO

[SG90] A. Schreiner and N. Geers. Supercomputer software from ODIN. Supercomputer, 7(2):81–88, March 1990. CODEN SPCOEL. ISSN 0168-7875.

Schonauer:1995:CIS

[SH95] W. Schönauer and H. Häfner. A careful interpretation of simple kernel benchmarks yields the essential information about a parallel supercomputer. Supercomputer, 11(4):63– 74, September 1995. CODEN SP-COEL. ISSN 0168-7875.

Simon:1997:SMP

[SH97] J. Simon and O. Heinz. SCI multiprocessor PC cluster in a Windows NT environment. Supercomputer, 13 (2):44–57, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Simon:1995:HPC

[Sim95] H. D. Simon. High performance computing in the US in 1994 — an analysis on the basis of the TOP500 list. Supercomputer, 11(2–3):21–30, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Simon:1997:TSS

[Sim97] H. D. Simon. The TOP25 supercomputer sites. Supercomputer, 13 (1):45–58, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

Shimizu:1991:SLR

[SK91] K. Shimizu and Y. Kanada. Solving linear recurrence problems on supercomputers. Supercomputer, 8 (1):30–37, January 1991. CODEN SPCOEL. ISSN 0168-7875.

Stewart:1994:IDR

[SNtW94] A. Stewart, M. Nool, H. J. J. te Riele, and D. T. Winter. An investigation of data reuse on the Cray S-MP System 500. Supercomputer, 11 (1):4–15, April 1994. CODEN SP-COEL. ISSN 0168-7875. Issue labeled April 1994 – April 1995.

Solchenbach:1989:ASS

[Sol89a] K. Solchenbach. Application software for Suprenum. Supercomputer, 6(2):44–50, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Solchenbach:1989:SFM

[Sol89b] K. Solchenbach. Suprenum-Fortran — an MIMD/SIMD language. Supercomputer, 6(2):25–30, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Solchenbach:1989:SMS

[Sol89c] K. Solchenbach. Suprenum-Fortran — an MIMD/SIMD language. Supercomputer, 6(2):25–30, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Synnott:1990:TDG

[SRS^{+}90] S. Synnott, J. Riedel, J. Stuve, P. Halamek, and W. Lehr. Threedimensional geometry from image processing on the JPL/CIT hypercube. Supercomputer, 7(1):21–28, January 1990. CODEN SPCOEL. ISSN 0168-7875.

Simon:1995:ALS

[SS95] H. D. Simon and E. Strohmaier. Amdahl’s Law and the statistical content of the NAS parallel benchmarks. Supercomputer, 11(4):75–88, September 1995. CODEN SPCOEL. ISSN 0168-7875.

Skourup:1988:VBI

[SSL88] J. Skourup, I. Svendsen, and J. Larsen. Vectorization of a boundary integral equation method. Supercomputer, 5(1):26–32, January 1988. CODEN SPCOEL. ISSN 0168-7875.

Streng:1996:PCM

[Sv96] M. Streng and R. van der Pas. Performance considerations for the MIPS R8000 microprocessor. Supercomputer, 12(3):4–20, August 1996. CODEN SPCOEL. ISSN 0168-7875.

Tedesco:1990:SNC

[TM90] J. Tedesco and P. McGill. Solution of natural convection in a 3-D cavity by the finite-element method. Supercomputer, 7(5):38–44, September 1990. CODEN SPCOEL. ISSN 0168-7875.

Thomas:1989:SCP

[TP89] B. Thomas and K. Peinze. Suprenum comfort of parallel programming. Supercomputer, 6(2):31–43, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Touzene:1993:PSA

[TP93] A. Touzene and B. Plateau. Parallel synchronous and asynchronous iterative methods to solve Markov chain problems. Supercomputer, 10(3):28– 39, May 1993. CODEN SPCOEL. ISSN 0168-7875.

Trottenberg:1989:SC

[Tro89] U. Trottenberg. Suprenum — the concept. Supercomputer, 6(2):5–12, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Tsuda:1991:DIV

[Tsu91] T. Tsuda. Design and implementation of a vectorizing compiler for the block-structured language Pascal. Supercomputer, 8(6):12–21, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Twerda:1996:PIT

[TVV96] A. Twerda, A. P. Van den Berg, and A. J. Van der Steen. Parallel implementation of time dependent Rayleigh-Benard convection. Supercomputer, 12(2):36–47, March 1996. CODEN SPCOEL. ISSN 0168-7875.

Uehara:1991:BVI

[UT91] T. Uehara and T. Tsuda. Benchmarking vector indirect load/store instructions. Supercomputer, 8(6): 57–74, November 1991. CODEN SP-COEL. ISSN 0168-7875.

VanderVorst:1988:VAS

[Van88a] H. Van der Vorst. Vectorial aspects of software libraries. Supercomputer, 5(1):33–41, January 1988. CODEN SPCOEL. ISSN 0168-7875.

Vorst:1988:VAS

[Van88b] H. A. Van der Vorst. Vectorization aspects of software. Supercomputer, 23(????):33–41, 1988. CODEN SP-COEL. ISSN 0168-7875.

VanderVorst:1989:EVC

[Van89] H. A. Van der Vorst. Eta-10P versus Cyber 205. Supercomputer, 6(1): 17–22, January 1989. CODEN SP-COEL. ISSN 0168-7875.

vanderSteen:1990:PPG

[van90a] A. van der Steen. Portable parallel generation of random numbers. Supercomputer, 7(1):18–20, January 1990. CODEN SPCOEL. ISSN 0168-7875.

VanderVorst:1990:EPV

[Van90b] H. A. Van der Vorst. Experiences with parallel vector computers for sparse linear systems. Supercomputer, 7(3):28–35, May 1990. CO-DEN SPCOEL. ISSN 0168-7875.

vanderSteen:1991:PAR

[van91] A. van der Steen. Performance analysis of radix-2 and radix-3 FFT algorithms for vector processors. Supercomputer, 8(3):9–13, May 1991. CODEN SPCOEL. ISSN 0168-7875.

VandenEynde:1992:SPB

[Van92a] D. Van den Eynde. Swell prediction off the Belgian coast — an operational application of supercomputing. Supercomputer, 9(2):37–45, March 1992. CODEN SPCOEL. ISSN 0168-7875.

vanderSteen:1992:PRS

[van92b] A. J. van der Steen. Preliminary results for some EuroBen V3.0 programs. Supercomputer, 9(6):41–46, November 1992. CODEN SPCOEL. ISSN 0168-7875.

vanderPas:1993:PIG

[van93a] R. van der Pas. The PVM implementation of a Generalized Red Black algorithm. Supercomputer, 10 (4–5):72–85, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

VanDerSteen:1993:EBR

[Van93b] A. J. Van Der Steen. EuroBen Benchmark results for the Hitachi S3800. Supercomputer, 10(4–5):32– 45, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

vanderSteen:1993:SDE

[van93c] A. J. van der Steen. Status and direction of the EuroBen Benchmark. Supercomputer, 10(4–5):19– 31, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

vanderSteen:1995:SDA

[Van95a] A. Van der Steen. Short description of architectures in the TOP500. Supercomputer, 11(2–3):61–93, June 1995. CODEN SPCOEL. ISSN 0168-7875.

VanderSteen:1995:SDE

[Van95b] A. J. Van der Steen. Status and direction of the EuroBen Benchmark. Supercomputer, 11(4):4–18, September 1995. CODEN SPCOEL. ISSN 0168-7875.

VanDerSteen:1996:SDN

[Van96] A. Van Der Steen. Short description of newly featuring architectures in the TOP500. Supercomputer, 12(1): 48–55, January 1996. CODEN SP-COEL. ISSN 0168-7875.

vanderSteen:1997:SDN

[Van97] A. Van der Steen. Short description of newly featuring architectures in the TOP500. Supercomputer, 13 (1):59–73, ???? 1997. CODEN SP-COEL. ISSN 0168-7875.

VanderLinden:1992:UEP

[VD92a] R. A. M. Van der Linden and A. M. De Meyer. User experiences with the parallel I/O access method. Supercomputer, 9(3):7–15, May 1992. CO-DEN SPCOEL. ISSN 0168-7875.

vanderSteen:1992:SRE

[vd92b] A. van der Steen and P. de Rijk. Summary of results of the EuroBen benchmark V2.0. Supercomputer, 9 (6):20–40, November 1992. CODEN SPCOEL. ISSN 0168-7875.

Vulk:1993:FEG

[Vul93] C. Vulk. Further experiences with GMRESR. Supercomputer, 10(3): 13–27, May 1993. CODEN SP-COEL. ISSN 0168-7875.

vanderSteen:1988:STC

[vv88] A. van der Steen and J. van Kats. Summary of tests on a Cyber 990, a Honeywell-Bull DPS-90, an IBM 3090, and a Unisys 1100/90 ISP. Supercomputer, 5(4):31–41, July 1988. CODEN SPCOEL. ISSN 0168-7875.

Wallace:1990:ST

[Wal90] D. J. Wallace. Supercomputing with transputers. Supercomputer, 7(2): 120–131, March 1990. CODEN SP-COEL. ISSN 0168-7875.

Waser:1993:PB

[Was93] S. Waser. The ParStone benchmark. Supercomputer, 10(4–5):46– 57, July–September 1993. CODEN SPCOEL. ISSN 0168-7875.

Werner:1989:SUI

[WBT89] K. H. Werner, U. Brass, and E. Thomas. The Suprenum user interface. Supercomputer, 6(2):20– 24, March 1989. CODEN SPCOEL. ISSN 0168-7875.

Wasniewski:1989:ILB

[WCMJ89] J. Wasniewski, J. Du Croz, P. Mayes, and L. Jankowski. Implementing the Level 2 BLAS on the Amdahl vector processors. Supercomputer, 6(5):34–43, September 1989. CODEN SPCOEL. ISSN 0168-7875.

Walker:1996:MSM

[WD96] D. W. Walker and J. J. Dongarra.

MPI: a standard Message Passing Interface. Supercomputer, 12(1):56– 68, January 1996. CODEN SP-COEL. ISSN 0168-7875.

Wang:1991:FSM

[WE91] Chia-Jiu Wang and M. A. Ess. A fast sparse-matrix solver for circuit simulation. Supercomputer, 8(3):22– 31, May 1991. CODEN SPCOEL. ISSN 0168-7875.

Wong:1991:SBP

[WGOY91] W. F. Wong, E. Goto, Y. Oyanagi, and N. Yoshida. Six benchmark problems for number crunchers. Supercomputer, 8(6):39, 42–43, November 1991. CODEN SPCOEL. ISSN 0168-7875.

Walde:1990:PFF

[WH90] W. Walde and O. Haan. Performance of Fast Fourier Transforms on vectorcomputers. Supercomputer, 7 (6):42–49, November 1990. CODEN SPCOEL. ISSN 0168-7875.

Witten:1991:SBM

[Wit91] M. Witten. Supercomputers in biology and medicine — an overview. Supercomputer, 8(2):37–53, March 1991. CODEN SPCOEL. ISSN 0168-7875.

Wilson:1991:CCE

[WM91] S. Wilson and D. Moncrieff. Concurrent computation in electron correlation calculations for molecules: ccMBPT. Supercomputer, 8(5):28– 40, September 1991. CODEN SP-COEL. ISSN 0168-7875.

Wu:1990:VMD

[WTW90a] Chwan-Hwa Wu, Jyun-Hwei Tsai, and Chia-Jiu Wang. Vectorized and multitasked 2-D simulations of RF discharges on a Cray X-MP. Supercomputer, 7(4):17–29, July 1990. CODEN SPCOEL. ISSN 0168-7875.

Wu:1990:VMS

[WTW90b] Chwan-Hwa Wu, Jyun-Hwei Tsai, and Chia-Jiu Wang. Vectorized and multitasked 2-D simulations of RF discharges on a Cray X-MP. Supercomputer, 7(4):17–29, July 1990. CODEN SPCOEL. ISSN 0168-7875.

Zhang:1989:PBS

[Zha89] Xiaodong Zhang. Parallel block SOR methods and various effects on shared and local memory multiprocessors. Supercomputer, 6(3):24–35, ???? 1989. CODEN SPCOEL. ISSN 0168-7875.

Zilli:1993:PIR

[Zil93] G. Zilli. Parallel implementation of a row-projection method for solving sparse linear systems. Supercomputer, 10(1):33–43, January 1993. CODEN SPCOEL. ISSN 0168-7875.

Zilli:1996:PMS

[Zil96] G. Zilli. Parallel method for sparse nonsymmetric linear and nonlinear systems of equations on a transputer network. Supercomputer, 12(4):4– 15, December 1996. CODEN SP-COEL. ISSN 0168-7875.

Zois:1987:PPF

[Zoi87] Demetris Zois. PARFES, a parallel finite element system. Supercomputer, 17 (??)(??):34–43, January 1987. CODEN SPCOEL. ISSN 0168-7875.

Zemerly:1995:CBB

[ZPN95] M. J. Zemerly, J. Papay, and G. R.

Nudd. Characterisation based bottleneck analysis of parallel systems. Supercomputer, 11(4):89–101, September 1995. CODEN SPCOEL. ISSN 0168-7875.

Zlatev:1992:OAP

[ZWMV92] Z. Zlatev, J. Wasniewski, J. Moth, and M. Venugopal. Optimizing air pollution models on parallel machines. Supercomputer, 9(2):23–36, March 1992. CODEN SPCOEL. ISSN 0168-7875.

Zlatev:1989:RCG

[ZWS89] Z. Zlatev, J. Wasniewski, and K. Schaumburg. Running conjugate gradient algorithms on three vector machines. Supercomputer, 6(1):

31–41, January 1989. CODEN SP-COEL. ISSN 0168-7875.